OpenStudy (anonymous):

The length of a rectangle is 3 inches more than its width. The area of the rectangle is 40 square inches. What is the length, in inches, of the rectangle? (1) 5 (3) 8.5 (2) 8 (4) 11.5

OpenStudy (chaise):

A=LxW A=(w+3)*w 40=(w+3)*w 40=w^2+3w w^2+3w-40=0 w=5 The width of the rectangle is equal to 5, the length (because the length is w+3) is equal to 8.

OpenStudy (anonymous):

2 is the answer

OpenStudy (kropot72):

I get the same answer as chaise: length = 8 inches